OpenStudy (anonymous):

Find all values of the parameter p for which the integral, 1/(x(lnx)^p) from e to infinity, converges.

OpenStudy (turingtest):

let \(u=\ln x\implies du=\frac{dx}x\) what do you get after integrating?

OpenStudy (anonymous):

the integral of 1/(u)^p du?

OpenStudy (turingtest):

yes, now integrate...

OpenStudy (anonymous):

u^(-p+1)/(-p+1)

OpenStudy (turingtest):

right, unless p=1, in which case you get a different case evaluate what you have for now

OpenStudy (turingtest):

think about three possibilities p>1 p=1 p<1

OpenStudy (turingtest):

think about this as\[\lim_{n\to\infty}\left.\frac1{p-1}\frac1{u^{p-1}}\right|_1^n\]

OpenStudy (turingtest):

\[=\lim_{n\to\infty}\frac1{p-1}\left(\frac1{n^{p-1}}-1\right)\]again consider three possibilities: p<1, P=1, p>1

OpenStudy (anonymous):

is p>1 because its 1/p-1?

OpenStudy (turingtest):

p=1 is a sort of breaking point for the cases for that reason, but also and more importantly because this limit will only converge if the exponent in the denominator is positive, right?

OpenStudy (turingtest):

oh I see what you are saying, so yes, you are right you still have to consider the case when p=1 though

OpenStudy (turingtest):

it's pretty trivial though

OpenStudy (turingtest):

I mean you should do the other integral though, just to show that it diverges as well

OpenStudy (turingtest):

...at least I would to be thorough, but as you prefer

OpenStudy (anonymous):

okay thank you!